OpenStudy (anonymous):

Divide. (8x^2-20+15) / (2x-3)

OpenStudy (rane):

use calculator

OpenStudy (anonymous):

don't have one :(

OpenStudy (anonymous):

mathway.com

OpenStudy (anonymous):

8x^2-5/2x-3

OpenStudy (anonymous):

I tried it just now it didn't give me an answer?

OpenStudy (anonymous):

well the answer is 8x^2 -5 all over 2x-3 do you need an explanation?

OpenStudy (jack1):

is that just a-20 or should it be -20x...?

OpenStudy (anonymous):

yes sorry it should be -20x

OpenStudy (jack1):

have you tried doing long division...?

OpenStudy (jack1):

for example, have a look at this:

OpenStudy (anonymous):

We can say that : \[8x^2-20x+3=(2x-3)(ax+b)+c\] Now, trying to find a , b and c. We have : \[(2x-3)(ax+b)+c=2ax^2+2bx-3ax-3b+c=2ax^2+(2b-3a)x+(-3b+c)\] So : \[2a=8\\2b-3a=-20\\-3b+c=3\] Then : \[a=4\\b=-4\\c=-9\]

OpenStudy (anonymous):

first it should be 20x and the Quotient is (4x+4) Remainder is 3